# Weak Quasistability and Rajchman measures

### v1

@@ 7. Rajchman Measures and Weak Quasistability
where $\mu_{sc}$ and $\mu_{sd}$ are the singular-continuous and singular-discrete (or simply discrete) parts of $\mu.$ This is the canonical decomposition of $\mu$, where any of the theses parts may be absent. The absolutely continuous and the singular-continuous parts make up the continuous part of $\mu$:
@@
Let $\mu$ be a finite measure on $\Re_{\mathbb{T}\kern 0.5pt}.$ The Wiener characterisation [20, Eq.(31) p.81] of continuous measures (see also, e.g., [14, p.365]) says that $\mu$ is continuous if and only if the following mean of bounded summands goes to zero:
@@
By ()(*) there is a subsequence $\{n_{j}\}$ for which $\lim_{j}\int_{\mathbb{T}\kern 0.5pt}z^{n_{j}}d\kern 1.0pt\eta_{g}=0$}, and so by ()(**)
@@
Remark $\kern-0.5pt7\kern-0.5pt.4.$ A weakly quasistable operator with a common weak stability subsequence. A weakly unstable operator that is weakly quasistable was exhibited in Example 4.1. Theorem 7.3 ensures the existence of a special class of such operators, namely,
@@
$\int_{\mathbb{T}\kern 0.5pt}\!z^{m_{j}}\,d\mu\not\to 0\;\;\hbox{as}\;\;j\to\infty\quad\hbox{\it but}\;\quad{\lim}_{\,j}\!\int_{\mathbb{T}\kern 0.5pt}\!z^{n_{j}}\,d\mu=0.$
@@
$U^{m_{j}}_{\varphi,\mu}\psi\kern-1.0pt{{{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}0\;\;\hbox{for some}\;\;\psi\in L^{2}({\mathbb{T}\kern 0.5pt},\mu)\quad\;\hbox{\it but}\;\quad U^{n_{j}}_{\varphi,\mu}\psi\kern-1.0pt{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}0\;\;\hbox{for every}\;\;\psi\in L^{2}({\mathbb{T}\kern 0.5pt},\mu).$

### v2

@@ 7. Rajchman Measures and Weak Quasistability
where $\mu_{sc}$ and $\mu_{sd}$ are the singular-continuous and singular-discrete (or simply discrete) parts of $\mu.$ This is the canonical decomposition of $\mu$, where any of theses parts may be absent. The continuous part $\mu_{c}$ of $\mu$ consists of the absolutely continuous and singular-continuous parts of $\mu$:
@@
Let $\mu$ be a finite measure on $\Re_{\mathbb{T}\kern 0.5pt}.$ The Wiener characterisation [21, Eq.(31) p.81] of continuous measures (see also, e.g., [15, p.365]) says that $\mu$ is continuous if and only if the following mean of bounded summands goes to zero:
@@
By ()(*), there is a subsequence $\{n_{j}\}$ for which $\lim_{j}\int_{\mathbb{T}\kern 0.5pt}z^{n_{j}}d\kern 1.0pt\eta_{g}=0$}, and so by ()(**)
@@
Remark $\kern-0.5pt7\kern-0.5pt.4.$ A weakly quasistable operator with a common weak stability subsequence.
A weakly unstable operator that is weakly quasistable was exhibited in Example 4.1. Theorem 7.3 ensures the existence of a special class of such operators, namely,
@@
$\int_{\mathbb{T}\kern 0.5pt}\!z^{m_{j}}\,d\mu\not\to 0\;\;\hbox{as}\;\;j\to\infty\qquad\!\hbox{\it but}\qquad{\lim}_{\,j}\!\int_{\mathbb{T}\kern 0.5pt}\!z^{n_{j}}\,d\mu=0.$
@@
$U^{m_{j}}_{\varphi,\mu}\psi\kern-1.0pt{{{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}0\;\;\hbox{for some}\;\;\psi\in L^{2}({\mathbb{T}\kern 0.5pt},\mu)\quad\;\hbox{\it but}\;\quad U^{n_{j}}_{\varphi,\mu}\psi\kern-1.0pt{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}0\;\;\hbox{for every}\;\;\psi\in L^{2}({\mathbb{T}\kern 0.5pt},\mu).$
Acknowledgement
I thank my co-authors in reference [9] for several discussions we had on Rajchman measures and weak stability.